Tremaine would like to organize UTA as either an S Corporation or a C corporation. In either form, the entity will generate a 10 percent annual before-tax return on a $1,000,000 investment. Tremaine’s marginal income tax rate is 37 percent and his tax rate on dividends and capital gains is 23.8 percent (including the net investment income tax). If Tremaine organizes UTA as an S corporation he will be allowed to claim the deduction for qualified business income. Also, because Tremaine will participate in UTA’s business activities, the income from UTA will not be subject to the net investment income tax. Assume that UTA will pay out 100 percent of its after-tax earnings every year as a dividend if it is formed as a C corporation. How much would cash after-taxes would Tremaine receive from his investment in the first year if UTA is organized as either an S corporation? How much would cash after-taxes would Tremaine receive from his investment in the first year if UTA is organized as either a C corporation?
Answer #1

B А 1 Particular 2 Pretax earnings 3 QBI deduction (20%) 4 Taxable income 5 Income tax (37%) 6 Cash after taxes s CorporationB s Corporation =1000000*10% =B2*20% 1 Particular 2 Pretax earnings 3 QBI deduction (20%) 4 Taxable income 5 Income tax (37%)